# Lagrangian Skeleta and Koszul Duality on Bionic Symplectic Varieties

### v1

@@ 3.3. Duality and Hamiltonian Reduction
We require a better understanding of the relation between $\widehat{A}$-modules and $A$-modules. Certainly, there is a flat morphism of algebras $A\to\widehat{A}$, and the corresponding functor $M\to\widehat{A}\otimes_{A}M$; which agrees with the naive completion when $M$ is finite over $A$. In certain cases we can can actually recover $M$ from its completion. To state the relevant result, recall the adjoint pair of functors $\mathbb{H}\colon A\text{-}{\mathsf{Mod}}\to B\text{-}{\mathsf{Mod}}$ and $\mathbb{H}^{\perp}\colon B\text{-}{\mathsf{Mod}}\to A\text{-}{\mathsf{Mod}}$ given by
@@ 3.4. Preservation of Holonomicity
${\;\stackrel{{{}_{\sim}}}{{\to}}\;}\mathbb{R}\operatorname{Hom}_{A_{\hbar}}(A_{\hbar}/J_{\hbar},A_{\hbar})\otimes_{A_{\hbar}}^{\mathbb{L}}M{\;\stackrel{{{}_{\sim}}}{{\to}}\;}J_{\hbar}^{\prime}\backslash A_{\hbar}\otimes_{A_{\hbar}}^{\mathbb{L}}[-n],$
where we applied the isomorphism (3.4). This establishes the first claim.
@@
proving the second claim. ∎
@@ 3.6. Extensions of Modules
Now we wish to specialize the situation: we suppose from here on in that $\mathfrak{X}$ is a bionic symplectic resolution of singularities. Thus we have a proper birational $\mathsf{T}$-equivariant morphism $\mu:\mathfrak{X}\to X:=\mathfrak{X}^{\operatorname{aff}}$, where $X$ is an affine symplectic variety, conic with respect to the elliptic action. Fix $r\in I$ and let $U(r)$ be an affine $\mathsf{T}$-stable open subset of $\mathfrak{X}$ containing the fixed point $y_{r}$. Then $U(r)$ contains both $\mathsf{\Lambda}_{r}$ and $C_{r}$.
@@
The weights of the Hamiltonian ${\mathbb{G}}_{m}$-action on $\Gamma(U(r),\delta_{\mathsf{\Lambda}_{r}})^{{\mathbb{G}}_{m}}$ are bounded above.
@@
By definition, $\delta_{\mathsf{\Lambda}_{r}}$ is supported on the attracting locus of the Hamiltonian ${\mathbb{G}}_{m}$-action on $U(r)$. Therefore, one can argue just as in the proof of Proposition 3.5. Namely, $M:=\Gamma(U(r),\delta_{\mathsf{\Lambda}_{i}})^{{\mathbb{G}}_{m}}$ has a good filtration whose associated graded is a $\Gamma(U(r),\mathcal{O})$-module supported on $\mathsf{\Lambda}_{r}$. It follows that the weights of $\operatorname{gr}M$ are bounded above. The same must therefore hold for $M$. ∎
@@
The ${\mathcal{W}}_{U(r)}$-module $\delta_{\mathsf{\Lambda}_{r}}$ admits an extension to a holonomic ${\mathcal{W}}_{\mathfrak{X}}$-module which is set-theoretically supported on $\mathsf{\Lambda}$.
@@
Since we have assumed that $\mathfrak{X}$ is a symplectic resolution, it is known [BLPW2] that ${\mathcal{W}}\text{-}{\mathsf{good}}$ is equivalent to ${\mathcal{W}}^{\prime}\text{-}{\mathsf{good}}$ where ${\mathcal{W}}^{\prime}$ is a quantization for which localization holds. Furthermore, this equivalence is given by twisting by a line bundle; so that the object $\delta_{\mathsf{\Lambda}_{i}}$ constructed above is sent to the corresponding module $\delta^{\prime}_{\mathsf{\Lambda}_{i}}$ over ${\mathcal{W}}^{\prime}$. In addition, it follows that this twisting preserves the support of a coherent module. Thus to prove the proposition we can assume that localization holds for ${\mathcal{W}}$.
@@
We denote the attracting set in $X$ with respect to the Hamiltonian action by $\mathfrak{n}$. Then it follows from the properness of $\mu$ that $\mu^{-1}(\mathfrak{n})=\mathsf{\Lambda}$. Moreover, we have shown that $\Gamma(\mathfrak{X},\mathscr{M})^{{\mathbb{G}}_{m}}$ is supported on $\mathfrak{n}$. But, applying the localization functor to $\Gamma(\mathfrak{X},\mathscr{M})^{{\mathbb{G}}_{m}}$, it follows that $\mathscr{M}$ is supported on $\mu^{-1}(\mathfrak{n})=\mathsf{\Lambda}$ and we are done. ∎
3.7. An example
Consider the Kleinian singularity $Y_{n}$ of type $A_{n}$ and its minimal resolution $X_{n}$, a smooth symplectic surface. (The symplectic form is induced from volume form on ${\mathbb{C}}^{2}$). There is a natural two-torus $\mathsf{T}$ acting on $X_{n}$, and it has $n$ fixed points $\{p_{1},p_{2},\ldots,p_{n}\}$ all lying in the exceptional fibre of the resolution: indeed if the exceptional locus is $D_{1}\cup D_{2}\ldots\cup D_{n-1}$ where each $D_{i}$ is a projective line then $D_{i}\cap D_{i+1}=\{p_{i+1}\}$ for $21\leq i\leq n-$ 2 and the other fixed points $p_{1}$ and $p_{n}$ lie on $D_{1}\backslash\{p_{1}\}$ and $D_{n-1}\backslash\{p_{n-1}\}$ respectively.
The Hamiltonian $\mathbb{G}_{m}$-action $\mathsf{H}\colon\mathbb{G}_{m}\to\mathsf{T}$ is induced from a Hamiltonian action on ${\mathbb{C}}^{2}$, and the locus of points $\mathsf{\Lambda}$ which have a limit under it is the union of the exceptional locus and an affine line tangent to $C_{n-1}$ at $p_{n}$. On the open cell, one of course already has a symplectic variety, so it is its own reduction, while on each of the other cells, the reduction is just a point.
Quantizations for $X_{n}$ have been explicitly studied in [Ku], and using his work one sees that for a generic choice of quantization, the simple extension $\mathcal{L}_{i}$ of the unique simple module supported on the Lagrangian $\mathsf{\Lambda}_{i}$ will be $\bigcup_{0\leq j\leq i}\overline{\mathsf{\Lambda}}_{j}$, and so the multiplicity of $\mathcal{L}=\bigoplus_{i=0}^{n}\mathcal{L}_{i}$ along $C_{i}$ is $n-i+1$. For special quantizations however, there may be different multiplicities. Note that one could chose other extensions of the simple module on $C_{i}$: for example the sheaves corresponding to standard modules in category $\mathcal{O}$ for the associated rational Cherednik algebra. In the generic case these coincide with the simple extensions, and their supports independent of the choice of equivariant quantization.
@@ 4. Koszul Duality and local generators
We begin by recalling that ${\mathcal{W}}\text{-}{\mathsf{good}}$ is the abelian category of ${\mathbb{G}}_{m}$-equivariant good ${\mathcal{W}}$-modules and we set ${\mathcal{W}}\text{-}{\mathsf{qcoh}}=\operatorname{Ind}{\mathcal{W}}\text{-}{\mathsf{good}}$. Here, and for the remainder of the article, ${\mathbb{G}}_{m}$ denotes the elliptic action. We will often take ${\mathbb{G}}_{m}$-invariants of an ${\mathbb{G}}_{m}$-equivariant sheaf $\mathscr{F}$. Let $\mathfrak{X}_{\mathrm{Eq}}$ denote the space $\mathfrak{X}$, equipped with the equivariant topology where the open sets are the ${\mathbb{G}}_{m}$-stable open sets in $\mathfrak{X}$. Then $\mathscr{F}^{{\mathbb{G}}_{m}}$ is the sheaf on $\mathfrak{X}_{\mathrm{Eq}}$ defined by $\mathscr{F}^{{\mathbb{G}}_{m}}(U)=\mathscr{F}(U)^{{\mathbb{G}}_{m}}$. Here $\beta\colon\mathfrak{X}\to\mathfrak{X}_{\mathrm{Eq}}$ is the map of topological spaces which is the identity on points.
@@ 4.1. $\Omega$-modules
    ${\mathcal{W}}\cong\mathcal{E}^{\vee}\otimes_{\Omega_{\hbar}}\mathcal{E}$ as dg-algebras, where the differential on ${\mathcal{W}}$ is trivial.
@@
Part (2). The evaluation map is a morphism of dg-algebras $\mathcal{E}^{\vee}\otimes_{\Omega_{\hbar}}\mathcal{E}\to{\mathcal{W}}$. To show that it is an isomorphism, it suffices to check locally that this is an isomorphism of sheaves (forgetting the grading and differential). In this case we may assume that $\mathcal{E}={\mathcal{W}}^{\oplus k}$ and the claim is clear.
@@
    $\mathcal{E}$ is not $K$-flat (in the sense of Spaltenstein) as left dg $\Omega_{\hbar}$-module.
@@
Identifying $C({\mathcal{W}}\text{-}{\mathsf{qcoh}})=\operatorname{Ind}C^{b}({\mathcal{W}}\text{-}{\mathsf{good}})$ and using the fact that $G$ and $F$ are cocontionous on $C^{b}({\mathcal{W}}\text{-}{\mathsf{good}})$ and $\Omega\text{-}{\mathsf{mod}}$ respectively, the above proof shows that $G$ and $F$ extend to equivalences $C({\mathcal{W}}\text{-}{\mathsf{qcoh}}){\;\stackrel{{{}_{\sim}}}{{\to}}\;}\Omega\text{-}{\mathsf{Mod}}$ and $\Omega\text{-}{\mathsf{Mod}}{\;\stackrel{{{}_{\sim}}}{{\to}}\;}C({\mathcal{W}}\text{-}{\mathsf{qcoh}})$ respectively. Passing to homotopy categories gives the final claim of the lemma. ∎
@@ 4.2. Existence of a local generator
$\mathcal{L}=\bigoplus_{r\in I}\mathbb{R}(j_{r})_{*}\delta_{\mathsf{\Lambda}_{r}}.$
@@
Next, we claim that $\{S_{i}:=\pi(C\cap U_{i})\}_{i\in I}$ is a filtered system of ${\mathbb{G}}_{m}$-stable affine open neighbourhoods of $s$ in $S$. It follows the explicit construction of $\pi$ given in Section 2.4 of [BDMN] that $S_{i}=\mathrm{Spec}((R/I)^{\{I,-\}})$, where $R=\mathcal{O}(U_{i})$ and $I$ is the radical ideal in $R$ defining $C$. Thus, each $S_{i}$ is an affine ${\mathbb{G}}_{m}$-stable open open neighbourhood of $s$. If $s\in U\subset S$ then there exists an open set $U^{\prime}\subset\mathfrak{X}$ containing $y$ such that $U^{\prime}\cap C=\pi^{-1}(U)$. If $i\in I$ such that $U_{i}\subset U^{\prime}$ then $S_{i}\subset U$.
@@
where we have applied the adjunction of Proposition 3.11 in the fourth line. ∎
@@ 4.3. The coderived category
Implicitly, the proof of Proposition 4.14 is based on classical $\mathcal{D}$-$\Omega$-duality, which is an equivalence $D^{b}(\mathcal{D}_{V}\text{-}{\mathsf{mod}}){\;\stackrel{{{}_{\sim}}}{{\to}}\;}D_{ex}(\Omega_{\mathsf{DR}}\text{-}{\mathsf{mod}})$, originally considered by Kapranov [Kap] and later Saito [Sa] and Beilinson-Drifeld [BD1, BD2]. Here $\Omega_{\mathsf{DR}}$ is the (sheaf) de Rham dg-algebra of poly-differential forms on $V$. More recently, Positselski [Po, Appendix B] has given a beautiful reformulation (and generalization) of this result in terms of coderived categories. Namely, one first considers the smallest thick triangulated subcategory $\text{Acyc}^{\text{co}}(\Omega_{\mathsf{DR}})$ of $K(\Omega_{\mathsf{DR}}\text{-}{\mathsf{Mod}})$ that is cocomplete and contains the total complex of every exact triple in $C(\Omega_{\mathsf{DR}})$. Then Positselski defines the coderived category $D^{\operatorname{co}}(\Omega_{\mathsf{DR}})$ as the Verdier quotient of $K(\Omega_{\mathsf{DR}}\text{-}{\mathsf{Mod}})$ by $\text{Acyc}^{\text{co}}(\Omega_{\mathsf{DR}})$. He proves that the $(\Omega_{\mathsf{DR}},\mathcal{D}_{V})$-bimodule $\mathcal{T}:=\Omega_{\mathsf{DR}}\otimes_{\mathcal{O}}\mathcal{D}_{V}$ (equipped with the de Rham differential) induces an equivalence
@@ 4.4. Symplectic resolutions
By Proposition 3.16, for each $1\leq r\leq N=|I|$, there exists a simple holonomic $\mathsf{T}$-equivariant (good) ${\mathcal{W}}$-module $\mathcal{L}_{r}$ that is supported on $\mathsf{\Lambda}$ and $\mathcal{L}_{r}|_{U(r)}\cong\delta_{\mathsf{\Lambda}_{r}}$. We take now
@@
If $\mathcal{E}$ is a locally free resolution of $\mathcal{L}$ then $\Omega=\underline{\mathcal{E}nd}_{{\mathcal{W}}}(\mathcal{E})^{{\mathbb{G}}_{m}}$ is cohomologically supported on $\mathsf{\Lambda}$.
@@
We just need to show that $\mathcal{L}$ is a local generator for $D^{b}({\mathcal{W}}\text{-}{\mathsf{good}})$. Assume $\mathscr{M}\in D^{b}({\mathcal{W}}\text{-}{\mathsf{good}})$ with $\mathbb{R}\mathcal{H}om_{{\mathcal{W}}}(\mathcal{L},\mathscr{M})$ acyclic. If $U$ is the open coisotropic cell then $\mathcal{L}|_{U}=\mathcal{L}_{1}|_{U}=\delta_{\mathsf{\Lambda}_{1}}$ is a local generator by Proposition 4.14. As in the proof of Theorem 4.12, we may therefore assume by induction that there is a closed coisotropic cell $C=C_{N}$ with complement $U$ such that $\mathscr{M}|_{U}\cong 0$. Again, $\mathscr{M}=\mathbb{H}^{\perp}(\mathscr{M}^{\prime})$ for some good ${\mathcal{W}}_{S}$-module $\mathscr{M}^{\prime}$. By Proposition 4.16, we are reduced to showing that $\mathbb{H}^{*}(\mathcal{L})$ is a local generator on $S$. Note that

### v2

@@ 3.3. Duality and Hamiltonian Reduction
We require a better understanding of the relation between $\widehat{A}$-modules and $A$-modules. Certainly, there is a flat morphism of algebras $A\to\widehat{A}$, and the corresponding functor $M\to\widehat{A}\otimes_{A}M$, which agrees with the naive completion when $M$ is finite over $A$. In certain cases we can can actually recover $M$ from its completion. To state the relevant result, recall the adjoint pair of functors $\mathbb{H}\colon A\text{-}{\mathsf{Mod}}\to B\text{-}{\mathsf{Mod}}$ and $\mathbb{H}^{\perp}\colon B\text{-}{\mathsf{Mod}}\to A\text{-}{\mathsf{Mod}}$ given by
@@ 3.4. Preservation of Holonomicity
${\;\stackrel{{{}_{\sim}}}{{\to}}\;}\mathbb{R}\operatorname{Hom}_{A_{\hbar}}(A_{\hbar}/J_{\hbar},A_{\hbar})\otimes_{A_{\hbar}}^{\mathbb{L}}M{\;\stackrel{{{}_{\sim}}}{{\to}}\;}J_{\hbar}^{\prime}\backslash A_{\hbar}\otimes_{A_{\hbar}}^{\mathbb{L}}M[-n],$
where we applied the isomorphism (3.4). This establishes the first isomorphism.
@@
proving the second isomorphism. ∎
@@ 3.6. Extensions of Modules
Now we wish to specialize the situation: we suppose from here on in that $\mathfrak{X}$ is a bionic symplectic resolution of singularities. Thus we have a proper birational $\mathsf{T}$-equivariant morphism $\mu:\mathfrak{X}\to X:=\mathfrak{X}^{\operatorname{aff}}$, where $X$ is an affine symplectic variety, conic with respect to the elliptic action. Fix $r\in I$ and let $U(r)$ be an affine $\mathsf{T}$-stable open subset of $\mathfrak{X}$ containing the fixed point $y_{r}$. Then $U(r)$ contains both $\Lambda^{c}_{r}$ and $C_{r}$. Let $\delta_{\Lambda^{c}_{r}}$ denote the holonomic simple module on $U(r)$ constructed in Proposition 3.5.
@@
The weights of the Hamiltonian ${\mathbb{G}}_{m}$-action on $\Gamma(U(r),\delta_{\Lambda^{c}_{r}})^{{\mathbb{G}}_{m}}$ are bounded above.
@@
By definition, $\delta_{\Lambda^{c}_{r}}$ is supported on the attracting locus of the Hamiltonian ${\mathbb{G}}_{m}$-action on $U(r)$. Therefore, one can argue just as in the proof of Proposition 3.5. Namely, $M:=\Gamma(U(r),\delta_{\Lambda^{c}_{i}})^{{\mathbb{G}}_{m}}$ has a good filtration whose associated graded is a $\Gamma(U(r),\mathcal{O})$-module supported on $\Lambda^{c}_{r}$. It follows that the weights of $\operatorname{gr}M$ are bounded above. The same must therefore hold for $M$. ∎
@@
The ${\mathcal{W}}_{U(r)}$-module $\delta_{\Lambda^{c}_{r}}$ admits an extension to a holonomic ${\mathcal{W}}_{\mathfrak{X}}$-module which is set-theoretically supported on $\Lambda^{c}$.
@@
Since we have assumed that $\mathfrak{X}$ is a symplectic resolution, it is known [BLPW2] that ${\mathcal{W}}\text{-}{\mathsf{good}}$ is equivalent to ${\mathcal{W}}^{\prime}\text{-}{\mathsf{good}}$ where ${\mathcal{W}}^{\prime}$ is a quantization for which localization holds. Furthermore, this equivalence is given by twisting by a line bundle; so that the object $\delta_{\Lambda^{c}_{i}}$ constructed above is sent to the corresponding module $\delta^{\prime}_{\Lambda^{c}_{i}}$ over ${\mathcal{W}}^{\prime}$. In addition, it follows that this twisting preserves the support of a coherent module. Thus to prove the proposition we can assume that localization holds for ${\mathcal{W}}$.
@@
We denote the attracting set in $X$ with respect to the Hamiltonian action by $\mathfrak{n}$. Then it follows from the properness of $\mu$ that $\mu^{-1}(\mathfrak{n})=\Lambda^{c}$. Moreover, we have shown that $\Gamma(\mathfrak{X},\mathscr{M})^{{\mathbb{G}}_{m}}$ is supported on $\mathfrak{n}$. But, applying the localization functor to $\Gamma(\mathfrak{X},\mathscr{M})^{{\mathbb{G}}_{m}}$, it follows that $\mathscr{M}$ is supported on $\mu^{-1}(\mathfrak{n})=\Lambda^{c}$ and we are done. ∎
Example 3.17.
Continue with the notation of 2.10. Quantizations for $X_{n}$ have been explicitly studied in [Ku], and using his work one sees that for a generic choice of quantization, the simple extension $\mathcal{L}_{i}$ of the unique simple module supported on the Lagrangian $\Lambda^{c}_{i}$ will be supported on $\bigcup_{0\leq j\leq i}\overline{\Lambda^{c}}_{j}$, and so the multiplicity of $\mathcal{L}=\bigoplus_{i=0}^{n}\mathcal{L}_{i}$ along $C_{i}$ is $n-i+1$. For special quantizations however, there may be different multiplicities. Note that one could chose other extensions of the simple module on $C_{i}$: for example the sheaves corresponding to standard modules in category $\mathcal{O}$ for the associated rational Cherednik algebra. In the generic case these coincide with the simple extensions, and their supports are independent of the choice of equivariant quantization.
@@ 4. Koszul Duality and local generators
[…]
@@ 4.1. $\Omega$-modules
    Suppose that the object $\mathcal{L}$ is nonzero. Then ${\mathcal{W}}\cong\mathcal{E}^{\vee}\otimes_{\Omega_{\hbar}}\mathcal{E}$ as dg-algebras, where the differential on ${\mathcal{W}}$ is trivial.
@@
Part (2). The evaluation map is a morphism of dg-algebras $\mathcal{E}^{\vee}\otimes_{\Omega_{\hbar}}\mathcal{E}\to{\mathcal{W}}$. To show that it is an isomorphism, it suffices to check locally that this is an isomorphism of sheaves (forgetting the grading and differential). In this case we may assume that $\mathcal{E}={\mathcal{W}}^{\oplus k}$ for some $k>0$ and the claim is clear.
@@
    $\mathcal{E}$ will not in general be $K$-flat (in the sense of Spaltenstein) as a left dg $\Omega_{\hbar}$-module.
@@
Identifying $C({\mathcal{W}}\text{-}{\mathsf{qcoh}})=\operatorname{Ind}C^{b}({\mathcal{W}}\text{-}{\mathsf{good}})$ and using the fact that $G$ and $F$ are cocontinuous on $C^{b}({\mathcal{W}}\text{-}{\mathsf{good}})$ and $\Omega\text{-}{\mathsf{mod}}$ respectively, the above proof shows that $G$ and $F$ extend to equivalences $C({\mathcal{W}}\text{-}{\mathsf{qcoh}}){\;\stackrel{{{}_{\sim}}}{{\to}}\;}\Omega\text{-}{\mathsf{Mod}}$ and $\Omega\text{-}{\mathsf{Mod}}{\;\stackrel{{{}_{\sim}}}{{\to}}\;}C({\mathcal{W}}\text{-}{\mathsf{qcoh}})$ respectively. Passing to homotopy categories gives the final claim of the lemma. ∎
@@ 4.2. Existence of a local generator
$\mathcal{L}=\bigoplus_{r\in I}\mathbb{R}(j_{r})_{*}\delta_{\Lambda^{c}_{r}}.$
@@
Next, we claim that $\{S_{i}:=\pi(C\cap U_{i})\}_{i\in I}$ is a filtered system of ${\mathbb{G}}_{m}$-stable affine open neighbourhoods of $s$ in $S$. It follows from the explicit construction of $\pi$ given in Section 2.4 of [BDMN] that $S_{i}=\mathrm{Spec}((R/I)^{\{I,-\}})$, where $R=\mathcal{O}(U_{i})$ and $I$ is the radical ideal in $R$ defining $C$. Thus, each $S_{i}$ is an affine ${\mathbb{G}}_{m}$-stable open open neighbourhood of $s$. If $s\in U\subset S$ then there exists an open set $U^{\prime}\subset\mathfrak{X}$ containing $y$ such that $U^{\prime}\cap C=\pi^{-1}(U)$. If $i\in I$ such that $U_{i}\subset U^{\prime}$ then $S_{i}\subset U$.
@@
where we have applied the adjunction of Proposition 3.11 in the fourth line.
@@ 4.3. The coderived category
Implicitly, the proof of Proposition 4.14 is based on classical $\mathcal{D}$-$\Omega$-duality, which is an equivalence $D^{b}(\mathcal{D}_{V}\text{-}{\mathsf{mod}}){\;\stackrel{{{}_{\sim}}}{{\to}}\;}D_{ex}(\Omega_{\mathsf{DR}}\text{-}{\mathsf{mod}})$, originally considered by Kapranov [Kap] and later Saito [Sa] and Beilinson-Drinfeld [BD1, BD2]. Here $\Omega_{\mathsf{DR}}$ is the (sheaf) de Rham dg-algebra of poly-differential forms on $V$. More recently, Positselski [Po, Appendix B] has given a beautiful reformulation (and generalization) of this result in terms of coderived categories. Namely, one first considers the smallest thick triangulated subcategory $\text{Acyc}^{\text{co}}(\Omega_{\mathsf{DR}})$ of $K(\Omega_{\mathsf{DR}}\text{-}{\mathsf{Mod}})$ that is cocomplete and contains the total complex of every exact triple in $C(\Omega_{\mathsf{DR}})$. Then Positselski defines the coderived category $D^{\operatorname{co}}(\Omega_{\mathsf{DR}})$ as the Verdier quotient of $K(\Omega_{\mathsf{DR}}\text{-}{\mathsf{Mod}})$ by $\text{Acyc}^{\text{co}}(\Omega_{\mathsf{DR}})$. He proves that the $(\Omega_{\mathsf{DR}},\mathcal{D}_{V})$-bimodule $\mathcal{T}:=\Omega_{\mathsf{DR}}\otimes_{\mathcal{O}}\mathcal{D}_{V}$ (equipped with the de Rham differential) induces an equivalence
@@ 4.4. Symplectic resolutions
By Proposition 3.16, for each $1\leq r\leq N=|I|$, there exists a simple holonomic $\mathsf{T}$-equivariant (good) ${\mathcal{W}}$-module $\mathcal{L}_{r}$ that is supported on $\Lambda^{c}$ and $\mathcal{L}_{r}|_{U(r)}\cong\delta_{\Lambda^{c}_{r}}$. We take now
@@
If $\mathcal{E}$ is a locally free resolution of $\mathcal{L}$ then $\Omega=\underline{\mathcal{E}nd}_{{\mathcal{W}}}(\mathcal{E})^{{\mathbb{G}}_{m}}$ is cohomologically supported on $\Lambda^{c}$.
@@
We just need to show that $\mathcal{L}$ is a local generator for $D^{b}({\mathcal{W}}\text{-}{\mathsf{good}})$. Assume $\mathscr{M}\in D^{b}({\mathcal{W}}\text{-}{\mathsf{good}})$ with $\mathbb{R}\mathcal{H}om_{{\mathcal{W}}}(\mathcal{L},\mathscr{M})$ acyclic. If $U$ is the open coisotropic cell then $\mathcal{L}|_{U}=\mathcal{L}_{1}|_{U}=\delta_{\Lambda^{c}_{1}}$ is a local generator by Proposition 4.14. As in the proof of Theorem 4.12, we may therefore assume by induction that there is a closed coisotropic cell $C=C_{N}$ with complement $U$ such that $\mathscr{M}|_{U}\cong 0$. Again, $\mathscr{M}=\mathbb{H}^{\perp}(\mathscr{M}^{\prime})$ for some good ${\mathcal{W}}_{S}$-module $\mathscr{M}^{\prime}$. By Proposition 4.16, we are reduced to showing that $\mathbb{H}^{*}(\mathcal{L})$ is a local generator on $S$. Note that